A carbon compod contains 12.8% carbon,2.1% hydrogen ,85.1% of bromine . The molecular weight of the compound is 187.9. calculate the molecular formula

Grade:11

1 Answers

Vikas TU
6 years ago
C compund= 12.8 % of atm wt of c(12)+ 2.1 % of atm wt of H(1)+85.1% of atm wt of br(80) 
For one,For C= 12.8/12=1.067 
H=2.1/1=2.1 
Be=85.1/80=1.067 
Separating with 1.067 to get basic nuclear ratio,1.067/1.067=12.1/1.067=21.067/1.067 
Weight =12+(2*1)+80=94 
Be that as it may, weight is given as 187.9 
So (CH2Br)×2 =(C2H4Br2).